# Large mechanical squeezing through synthetic magnetism in optomechanics

### v1

@@ III Synthetic magnetism enhanced mechanical squeezing
To point out the enhancement of the mechanical squeezing in our proposed system, we need to derive the position variance of the nonlinear mechanical resonator. For this purpose, we must evaluate the covariance matrix whose elements $V_{ij}$ fulfill $V_{ij}=\frac{\langle u_{i}u_{j}+u_{j}u_{i}\rangle}{2}$. These matrix elements also satisfy the motional equation,
@@
with the diagonal diffusion matrix $D=Diag[\frac{\kappa}{2},\frac{\kappa}{2},\frac{\gamma_{a}}{2},\frac{\gamma_{a}}{2},\\ \frac{\gamma_{m}}{2}e^{2r}(2n_{th}+1),\frac{\gamma_{m}}{2}e^{-2r}(2n_{th}+1)]$, and the matrix $\rm M$ must meet the Routh-Huritz stability criterion where all its eigenvalues should have negative real parts. We have checked that our used parameters satisfy this stability criterion. As we aim to study the steady-state behavior of the mechanical squeezing, we use the long time limit of Equation 14 which is known as Lyaponuv equation,
@@
Figure 2: Mechanical position variance from Equation 17 versus $G_{a}$ and $\Delta_{a}$. The mechanical coupling strengh is $J_{m}=0$ for (a) and $J_{m}=0.1\omega_{m}$ for (b). The parameters used are $\omega_{m}/2\pi=1\rm{MHz}$, $g_{m}=10^{-4}\omega_{m}$, $\kappa=0.02\omega_{m}$, $\gamma_{a}=0.4\omega_{m}$, $\gamma_{m}=10^{-4}\omega_{m}$, $\eta=10^{-4}\omega_{m}$, $n_{th}=100$, $G_{m}=0.15\omega_{m}$, $\tilde{\Delta}=-\omega^{\prime}_{m}$, and $\theta=\pi/2$.
From the $V_{ij}$ elements, we can express the position variance ($\langle\delta q_{m}^{2}\rangle$) and the phonon number ($n_{eff}^{m}$) of the mechanical resonator as,
@@
where $\langle\delta q_{ZPF}^{2}\rangle=\frac{1}{2}$ is the zero-point fluctuations of the mechanical resonator. Before pointing out the effect of the mechanical coupling $J_{m}$ and its phase $\theta$ on the variance squeezing derived in Equation 17, we need to check the optimal acoustic resonance $\Delta_{a}^{\rm{opt}}$ corresponding to an efficient supression of the heating term $G_{m}^{{}^{\prime}}(\delta a_{1}^{\dagger}\delta b_{m}^{s\dagger}+\delta a_{1}\delta b_{m}^{s})$. For this purpose, we work at the red-sideband detuning ($\tilde{\Delta}=-\omega_{m}^{{}^{\prime}}$), and plot the variance by simultaneously sweepping the acoustic effective coupling $G_{a}$ and detuning $\Delta_{a}$ as shown in Figure 2. It can be seen that the largest squeezing is generated at the acoustic resonance $\Delta_{a}^{\rm{opt}}\approx 3.5\omega_{m}$ as in Chen et al. (2023). From now on, we fix $\Delta_{a}=\Delta_{a}^{\rm{opt}}$ from now on, and keep the red-sideband detuning resonance condition. In Figure 2(a), there is no squeezing generated near $G_{a}\sim 0$, or it is very weak. However, Figure 2(b) shows a large degree of squeezing generated even for $G_{a}=0$ around the optimal acoustic resonance $\Delta_{a}^{\rm{opt}}$. This feature reveals how the synthetic jauge induces the squeezing in our proposal even if the BSBS effect is not supported in the system. Moreover, the colorbars in these two figures display how the squeezed generated for $J_{m}\neq 0$ is stronger compared to when there is no mechanical coupling $J_{m}=0$. To further get inslight into the enhancement of the squeezing through the synthetic jauge, we represent the mechanical position variance together with its effective phonon number, both versus $J_{m}$ and $\theta$, as displayed in Figure 3. The corresponding mechanical effective phonon number $n_{eff}^{m}$ can be equally expressed in term of the $V_{ij}$ elements and it yields,
@@
Figure 3: (a) Mechanical position variance (Equation 17) and (b) mechanical effective phonon number (Equation 18) versus $J_{m}$ and $\theta$. Optimal squeezing parameters deduced from Figure 2 are used, $G_{m}=0.15\omega_{m}$, $G_{a}=0.124\omega_{m}$, and $\Delta_{a}^{\rm{opt}}$. The others parameters are those in Figure 2.
@@
It can be seen from Figure 3(a,b) that both variance and phonon number are modulated along the $\theta$ direction. For $J_{m}\leq 5\times 10^{-2}\omega_{m}$, the effect of the phase $\theta$ is not pronounced in these figures. For $J_{m}>5\times 10^{-2}\omega_{m}$, however, the variance (Figure 3a) exhibits sort of peaks at $\theta\equiv(n+\frac{1}{2})\pi$ and dips for $\theta\equiv n\pi$, with $n$ being an interger. This feature shows that significant squeezing is generated at $\theta\equiv(n+\frac{1}{2})\pi$, which reversely corresponds to an effective minimum phonon number as expected in Figure 3b (see near $J_{m}\sim 0.1\omega_{m}$ for instance). By paying attention to the colorbars of these figures, it can be figured out that the more the mechanical coupling $J_{m}$ increases, the better we cool down the mechanical resonator and that results to a stronger degree of squeezing. These figures point out the crucial role played by both $J_{m}$ and $\theta$ regarding the squeezing enhancement in our proposal. To further reveal the oscillatory feature of the squeezing above discussed, we have extracted the variance and phonon number from Figure 3 at $J_{m}\sim 0.1\omega_{m}$, and that is displayed in Appendix A.
@@
Figure 4: Left and right column depicts the same quantities, which are plotted for $J_{m}=0$ and $J_{m}=0.1\omega_{m}$, respectively. (a,b) display the position variance of the mechanical resonator over the Brillouin acoustic decay rate $\gamma_{a}$ for three specific couples of points ($G_{a},G_{m}$): ($0.05\omega_{m},0.08\omega_{m}$), ($0.15\omega_{m},0.124\omega_{m}$), and ($0.25\omega_{m},0.154\omega_{m}$). (c,d) analyze the robustness of the mechanical position variance versus the thermal noise $n_{th}$. (e,f) figure out the behavior of the mechanical position variance versus the Duffing coefficient $\eta$. The rest of the parameters are the same as in Figure 2.
In what follow, we aim to figure out the magnitude of the position variance in order to appreciate the degree of the squeezing enhancement induced by the synthetic jauge. By doing so, it can be easy to analyze the behavior of the variance over some parameters instead of just observing its related colorbar as in Figure 2 and Figure 3. Therefore, we pick up specific parameters in the ($G_{m},G_{a}$) space, and represent the variance over some interesting variables as displayed in Figure 4. The left and right columns in Figure 4 depict the mechanical position variance plotted over the same variable for $J_{m}=0$ and $J_{m}=0.1\omega_{m}$, respectively. The first interesting parameter considered here is the Brillouin acoustic decay rate $\gamma_{a}$. The BSBS process involved in our proposal relies on the fact that $\gamma_{a}\gg\kappa(\gg\gamma_{m})$. Therefore, it is crucial to point out the impact of this condition on the squeezing generation as shown in Figure 4(a,b). Without the mechanical coupling ($J_{m}=0$), Figure 4a shows how the variance grows up rapidly and reaches a saturation limit above $\gamma_{a}\gg\kappa=2\times 10^{-2}\omega_{m}$ as expected. Above this BSBS condition, the position variance settles on a sort of plateau, where the acoustic decay rate does no longer affect the behavior of the generated squeezing. This regime could be of great interests for quantum technologies involving squeezing such as quantum information processing, quantum sensing/metrology, and quantum computing. By considering the synthetic jauge ($J_{m}\neq 0$), Figure 4b shows a great enhancement of the degree of squeezing (from 00 to near $\rm{5dB}$) compared to when $J_{m}=0$ (compare Figure 4a to Figure 4b). Despite the fact that the position variance behaves with almost the same shape in both Figure 4(a,b), it is worth to highlight that the merit of the synthetic jauge has been to push the limit of the generated squeezing beyond the $3dB$. Furthermore, these figures show that the more the couplings ($G_{m},G_{a}$) are enhanced, better is the degree of the squeezing (see Appendix A for a large view).
@@
In Figure 4c, we have plotted the position variance over the thermal phonon mechanical excitation. The shadow erea depicts the region below the $\rm{3dB}$. It can be seen that when $G_{a}=0$, there is only a short window where the squeezing is above the $\rm{3dB}$ (full line) compared to when $G_{a}\neq 0$ (dashed line). This feature reveals the key role played by the BSBS effect, which supresses heating processes in the system Chen et al. (2023). This BSBS effect is further reinforced through the gap between the two lines in Figure 4c, showing how $G_{a}\neq 0$ has improved the degree of squeezing compared to the case $G_{a}=0$. By taking into account the synthetic jauge, the first observation in Figure 4d is the gap between the cases $G_{a}=0$ and $G_{a}\neq 0$ that has been efficiently reduced. Indeed, the squeezing generated for $G_{a}=0$ follows almost the one generated for $G_{a}\neq 0$, and they stay longer beyond the $\rm{3dB}$ compared to what is shown in Figure 4d. This confirms that the synthetic jauge contributes to suppress heating channels, inducing strong squeezing even for $G_{a}=0$, as ealier discussed in Figure 2. Furthermore, it can be observed that the squeezing resists more to the mechanical thermal noise when the synthetic jauge is accounted. In Figure 4(e,f), we display the mechanical position variance versus the Duffing nonlinear coefficient $\eta$. These figures show how the variance sharply grows for weak values of $\eta$, and settles quickly to a plateau like a saturation limit. However, it can be observed that the synthetic jauge brigdes the gap between the cases $G_{a}=0$ and $G_{a}\neq 0$ as previouly discussed. Moreover, Figure 4e shows how the degree of the generated squeezing for $G_{a}=0$ is below the $\rm{3dB}$, while it exceeds this limit for $J_{m}\neq 0$ as depicted in Figure 4f. Once again, this reveals the merit of the synthetic jauge that induces a strong squeezing even for $G_{a}=0$ in our proposal as shown in Figure 2.
@@
In order to have a concise representation of these squeezed states generated in our proposal, we move to the phase space where we provide Wigner function distribution for specific parameters. Owing to the Gaussian nature of the quantum noise, our linearized system can be captured through a single Gaussian Wigner function in the steady state defined as Banerjee et al. (2023)
@@
where $u_{m}=(\delta q_{m},\delta p_{m})^{T}$ is the column vector of the mechanical fluctuations and $V_{m}$ stands for the covariance matrix for the mechanical mode. The left column in Figure 5 depicts the Wigner function distribution of the mechanical resonator (Figure 5a), the mechanical position variance (Figure 5b) and the mechanical momentum variance (Figure 5c) for the synthetic modulation phase of $\theta=0$. Similarly, the right column displays the same quantities at $\theta=\frac{\pi}{2}$ i.e., the Wigner function distribution (Figure 5d), the position variance (Figure 5e) and the momentum variance (Figure 5f). These figures have been captured for a long time enough that the system has reached its steady state dynamic. Moreover, these figures have been extracted from Figure 3a for the mechanical coupling strengh fixed at $J_{m}=0.1\omega_{m}$. In Figure 5(a,d), the contour represented by a dark circle depicts the related coherent state, while the shrunken and expanded shape bounded by the white ellipsis features the corresponding generated squeezed state. As predicted from Figure 3a, the wigner function at $\theta=0$ (Figure 5a) shows less squeezing as compared to when $\theta=\pi/2$ (Figure 5d). Moreover, these Wigner function distributions confirm the squeezing along the position direction, while the quantum fluctuation is transfered to the momentum direction which is amplified. To further analyze these observations, we have plotted the dyanmical time-evolution of the involved variances. As expected, both position and momentum dynamical evolution at $\theta=0$ clearly show no squeezing (Figure 5(b,c)), which is revealed by an oscillatory behavior above the Standard Quantum Limit (SQL). However, we can qualitatively observe that there is less amount of noise along the position direction compared to the momentum direction as predicted by the corresponding Wigner distribution (Figure 5a). For $\theta=\pi/2$, however, Figure 5e shows squeezing degree below the SQL, while its corresponding momentum is far above the SQL, revealing that quantum fluctuation has been amplified along this direction as predicted from Figure 5d. From the above analysis, it can be seen that our Wigner function distribution together with the dynamical evolution of the variances agree well with the synthetic magnetism squeezing enhancement pointed out in Figure 3. This work provides an efficient scheme toward squeezing enhancement beyond $\rm{3dB}$ in optomechanical system that is based on backward stimulated Brillouin scattering effect. The generated squeezing under this scheme is robust enough against thermal noise compared to the case without the synthetic magnetism. This investigation can be extended to electromechanical systems, hybrid opto-electromechanical system and to connexe fields.

### v2

@@ III Synthetic magnetism enhanced mechanical squeezing
To illustrate the enhancement of the mechanical squeezing in our proposed scheme, we need to derive the position variance of the nonlinear mechanical resonator. For this purpose, we evaluate the covariance matrix, $V_{ij}$, where $V_{ij}=\frac{\langle u_{i}u_{j}+u_{j}u_{i}\rangle}{2}$. The covariance matrix satisfies the equation of motion,
@@
with the diagonal diffusion matrix $D=Diag[\frac{\kappa}{2},\frac{\kappa}{2},\frac{\gamma_{a}}{2},\frac{\gamma_{a}}{2},\\ \frac{\gamma_{m}}{2}e^{2r}(2n_{th}+1),\frac{\gamma_{m}}{2}e^{-2r}(2n_{th}+1)]$, and the matrix $\rm M$ must meet the Routh-Hurwitz stability criterion where all its eigenvalues should have negative real parts. The parameters are chosen in such a way that this stability criterion is met and they are experimentally feasible. As we aim to study the steady-state behavior of the mechanical squeezing, we use the long time limit of Equation 14 which is known as the Lyapunov equation,
@@
Figure 2: Mechanical position variance from Equation 17 versus $G_{a}$ and $\Delta_{a}$. The mechanical coupling strength is $J_{m}=0$ for (a) and $J_{m}=0.1\omega_{m}$ for (b). The parameters used are $\omega_{m}/2\pi=1\rm{MHz}$, $g_{m}=10^{-4}\omega_{m}$, $\kappa=0.02\omega_{m}$, $\gamma_{a}=0.4\omega_{m}$, $\gamma_{m}=10^{-4}\omega_{m}$, $\eta=10^{-4}\omega_{m}$, $n_{th}=100$, $G_{m}=0.15\omega_{m}$, $\tilde{\Delta}=-\omega^{\prime}_{m}$, and $\theta=\pi/2$.
From the $V_{ij}$ elements, we can obtain the position variance ($\langle\delta q_{m}^{2}\rangle$) of the mechanical resonator as follows:
@@
where $\langle\delta q_{ZPF}^{2}\rangle=\frac{1}{2}$ is the zero-point fluctuations of the mechanical resonator. One can clearly see that the position quadrature is getting squeezed. In order to effectively suppress the thermal effects at the the optimal acoustic resonance $\Delta_{a}^{\rm{opt}}$, as depicted in the terms, $G_{m}^{{}^{\prime}}(\delta a_{1}^{\dagger}\delta b_{m}^{s\dagger}+\delta a_{1}\delta b_{m}^{s})$, we work in the red-detuned regime ($\tilde{\Delta}=-\omega_{m}^{{}^{\prime}}$). The main parameters used for our numerical simulations are $\omega_{m}/2\pi=1\rm{MHz}$, $g_{m}=10^{-4}\omega_{m}$, $\kappa=0.02\omega_{m}$, $\gamma_{a}=0.4\omega_{m}$, $\gamma_{m}=10^{-4}\omega_{m}$, $\eta=10^{-4}\omega_{m}$, $n_{th}=100$, $G_{m}=0.15\omega_{m}$, and $\tilde{\Delta}=-\omega^{\prime}_{m}$. The other parameters such as $J_{m}$, $\theta$, $G_{a}$ and $\Delta_{a}$ are adjusted as latter on indicated on the related figures. Figure 2 exhibits the position variance as a function of the the acoustic effective coupling $G_{a}$ and detuning $\Delta_{a}$. It can be observed that the largest squeezing is generated at the acoustic resonance $\Delta_{a}^{\rm{opt}}\approx 3.5\omega_{m}$ and $G_{a}=0.124\omega_{m}$. For the rest of the work, we take $\Delta_{a}=\Delta_{a}^{\rm{opt}}$ and assume the red-sideband detuning resonance condition. It is revealed in Fig. 2(a) that no squeezing is generated near $G_{a}\sim 0$, or it is very weak. However, Fig.2(b) shows a large degree of squeezing generated even for $G_{a}=0$ around the optimal acoustic resonance $\Delta_{a}^{\rm{opt}}$. This feature reveals how the synthetic gauge induces the squeezing even if the BSBS effect is not supported in the system. Moreover, it could be seen that the squeezing generated in the presence of $J_{m}\neq 0$ is stronger compared to the case when there is no mechanical coupling $J_{m}=0$. To get further insight into the enhancement of the squeezing through the synthetic gauge, we study the mechanical position variance against the variation of both $J_{m}$ and $\theta$, as displayed in Fig.3(a). On the other hand, corresponding study for the effective phonon number is depicted in Fig.3(b). It should be noted that the effective phonon number corresponding to the mechanical oscillator $n_{eff}^{m}$ can also be expressed in term of the $V_{ij}$ elements, and it yields,
@@
Figure 3: (a) Mechanical position variance (Equation 17) and (b) effective (mechanical) phonon number (Equation 18) versus $J_{m}$ and $\theta$. Optimal squeezing parameters deduced from Figure 2 are used: $G_{m}=0.15\omega_{m}$, $G_{a}=0.124\omega_{m}$, and $\Delta_{a}^{\rm{opt}}$. The others parameters are similar to the ones in Figure 2.
@@
It can be seen from Figure 3(a,b) that both variance and phonon number are modulated along the $\theta$ direction. For $J_{m}\leq 5\times 10^{-2}\omega_{m}$, the effect of the phase $\theta$ is negligible. However, for $J_{m}>5\times 10^{-2}\omega_{m}$, the variance (Figure 3a) exhibits peaks at $\theta\equiv(n+\frac{1}{2})\pi$ and dips for $\theta\equiv n\pi$, $n$ being an integer. This feature shows that significant squeezing is generated at $\theta\equiv(n+\frac{1}{2})\pi$. This also corresponds to an effective minimum phonon number as expected in Figure 3b (see near $J_{m}\sim 0.1\omega_{m}$ for instance). By paying attention to the colorbars of these figures, it can be figured out that with an increase in mechanical coupling $J_{m}$, the mechanical resonator is getting cooled down significantly resulting in stronger degree of squeezing. These figures point out the crucial role played by both $J_{m}$ and $\theta$ regarding enhancement of squeezing in our proposal. To further reveal the oscillatory feature of the squeezing above discussed, we have extracted the position variance and phonon number from Figure 3 at $J_{m}\sim 0.1\omega_{m}$, and that is displayed in Appendix B.
@@
Figure 4: Left and right column depicts the same quantities, which are plotted for $J_{m}=0$ and $J_{m}=0.1\omega_{m}$, respectively. (a,b) display the position variance of the mechanical resonator over the Brillouin acoustic decay rate $\gamma_{a}$ for two specific couples of points ($G_{a},G_{m}$): ($0.05\omega_{m},0.08\omega_{m}$) for the full line, and ($0.15\omega_{m},0.124\omega_{m}$) for the dashed line. (c,d) analyze the robustness of the mechanical position variance versus the thermal noise $n_{th}$. (e,f) figure out the behavior of the mechanical position variance versus the Duffing coefficient $\eta$. In (c-f), the full line is for $G_{a}=0$ and the dashed line is for $G_{a}=0.124\omega_{m}$. The rest of the parameters are the same as in Figure 2.
@@
Next, we investigate the role of synthetic gauge in enhancing the squeezing of the position quadrature. We take specific parameters in the ($G_{m},G_{a}$) space, and in Fig. 4 depict how the position variance varies with $\gamma_{a}$,$n_{th}$ and $\eta$. The left and right columns in Figure 4 depict the mechanical position variance plotted over the same variable for $J_{m}=0$ and $J_{m}=0.1\omega_{m}$, respectively. The first interesting parameter considered here is the Brillouin acoustic decay rate $\gamma_{a}$. The BSBS process involved in our proposal relies on the fact that $\gamma_{a}\gg\kappa(\gg\gamma_{m})$. Therefore, it is crucial to point out the impact of this condition on the squeezing generation as shown in Figure 4(a,b). Without the mechanical coupling ($J_{m}=0$), Figure 4a shows how the variance increases rapidly and reaches a saturation limit above $\gamma_{a}\gg\kappa=2\times 10^{-2}\omega_{m}$ as expected. Above this BSBS condition, the position variance settles on a sort of plateau, where the acoustic decay rate does no longer affect the behavior of the generated squeezing. This regime could be of great interests for quantum technologies involving squeezing such as quantum information processing, quantum sensing and metrology, and quantum computing. By considering the synthetic gauge ($J_{m}\neq 0$), Figure 4b shows a great enhancement of the degree of squeezing (from 00 to near $\rm{5dB}$) compared to when $J_{m}=0$ (compare Figure 4a to Figure 4b). Despite the fact that the position variance behaves with almost the same shape in both Figure 4(a,b), it is worth to mention that the merit of the synthetic gauge has been to push the limit of the generated squeezing beyond the $\rm{3dB}$. Furthermore, these figures show that the more the couplings ($G_{m},G_{a}$) are enhanced, better is the degree of the squeezing (see Appendix B for a large view). In Figure 4c, we have plotted the position variance over the thermal phonon mechanical excitation. The shadow area depicts the region below the $\rm{3dB}$. It can be seen that when $G_{a}=0$, there is only a short window where the squeezing is above the $\rm{3dB}$ (full line) compared to when $G_{a}\neq 0$ (dashed line). This feature reveals the key role played by the BSBS effect, which suppresses the heating processes in the system Shen et al. (2021). This BSBS effect is further reinforced through the gap between the two lines in Figure 4c, showing how $G_{a}\neq 0$ has improved the degree of squeezing compared to the case $G_{a}=0$. By taking into account the synthetic gauge, we observe in Figure 4d that the gap between the cases $G_{a}=0$ and $G_{a}\neq 0$ has been efficiently reduced. Indeed, the squeezing generated for $G_{a}=0$ follows almost the one generated for $G_{a}\neq 0$, and they stay longer beyond the $\rm{3dB}$ compared to what is shown in Figure 4d. This confirms that the synthetic gauge contributes to suppress the heating channels, inducing strong squeezing even for $G_{a}=0$, as earlier discussed in Figure 2. Furthermore, it can be observed that squeezing is somehow less prone to the affect of thermal phonons in the presence of synthetic gauge. In Figure 4(e,f), we display the mechanical position variance versus the Duffing nonlinear coefficient $\eta$. These figures show how the variance sharply increases for weak values of $\eta$, and settles quickly to a plateau like a saturation limit. However, it can be observed that the synthetic gauge bridges the gap between the cases $G_{a}=0$ and $G_{a}\neq 0$ as previously discussed. Moreover, Figure 4e shows how the degree of the generated squeezing for $G_{a}=0$ is below the $\rm{3dB}$, while it exceeds this limit for $J_{m}\neq 0$ as depicted in Figure 4f. Once again, this highlights the merit of the synthetic gauge that induces a strong squeezing even for $G_{a}=0$ in our proposal as shown in Figure 2.
@@
In Fig. 5(a) and (d) we exhibit the Wigner function distribution for the squeezed state generated for some specific parameters. Owing to the Gaussian nature of the quantum noise, our linearized system can be described through a single Gaussian Wigner function in the steady state defined as Banerjee et al. (2023)
@@
where $u_{m}=(\delta q_{m},\delta p_{m})^{T}$ is the column vector of the mechanical fluctuations and $V_{m}$ stands for the covariance matrix for the mechanical mode. The left column in Figure 5 depicts the Wigner function distribution of the mechanical resonator (Figure 5a), the mechanical position variance (Figure 5b) and the mechanical momentum variance (Figure 5c) for the synthetic modulation phase of $\theta=0$. Similarly, the right column displays the same quantities at $\theta=\frac{\pi}{2}$ i.e., the Wigner function distribution (Figure 5d), the position variance (Figure 5e) and the momentum variance (Figure 5f). These figures have been plotted in the long time limit so that the system has reached its steady state. Moreover, these figures have been extracted from Figure 3a for the mechanical coupling strength fixed at $J_{m}=0.1\omega_{m}$. In Figure 5(a,d), the contour represented by a dark circle depicts the related coherent state, while the shrunken and expanded shape bounded by the white ellipse features the corresponding generated squeezed state. As predicted from Figure 3a, the Wigner function at $\theta=0$ (Figure 5a) shows less squeezing as compared to when $\theta=\pi/2$ (Figure 5d). Moreover, these Wigner function distributions confirm the squeezing along the position direction, while the quantum fluctuation is transferred to the momentum direction which is amplified. To further analyze these observations, we have plotted the dynamical time-evolution of the involved variances. As expected, both position and momentum dynamical evolution at $\theta=0$ clearly show no squeezing (Figure 5(b,c)), which is revealed by an oscillatory behavior above the Standard Quantum Limit (SQL). However, we can qualitatively observe that there is less amount of noise along the position direction compared to the momentum direction as predicted by the corresponding Wigner distribution (Figure 5a). For $\theta=\pi/2$, however, Figure 5e shows squeezing in position quadrature below the SQL, while its corresponding momentum is far above the SQL, revealing that quantum fluctuation has been amplified along this direction as predicted from Figure 5d. From the above analysis, it can be seen that our Wigner function distribution together with the dynamical evolution of the variances agree well with the synthetic magnetism squeezing enhancement pointed out in Figure 3. This work provides an efficient scheme towards enhancement of squeezing beyond $\rm{3dB}$ in optomechanical system that is based on the backward stimulated Brillouin scattering effect. The generated squeezing under this scheme is robust enough against thermal noise compared to the case without the synthetic magnetism. This investigation can be extended to electromechanical systems and hybrid opto-electromechanical systems.